# How Far Are We on the Decision-Making of LLMs? Evaluating LLMs’ Gaming Ability in Multi-Agent Environments

### v3

@@ 3.1 Cooperative Games
The vanilla setting for this game is $MIN=0$, $MAX=100$, and $R=\frac{2}{3}$. We show the choices made by all agents as well as the average and the winning numbers in Fig. 2(1). Key observations are: (1) In the first round, agents consistently select $50$ (or close to $50$), corresponding to the mean of a uniform distribution ranging from 00 to $100$. This behavior suggests that the model fails to recognize that the winning number is $\frac{2}{3}$ of the average. (2) As rounds progress, the average number selected decreases noticeably, demonstrating that agents are capable of adapting based on historical outcomes. Since the optimal strategy is to choose the $MIN$, the score in this game is given by $S_{1}=\frac{1}{NK}\sum_{ij}(C_{ij}-MIN)$, where $C_{ij}$ is the chosen number of player $i$ in round $j$. The model scores [^2] For clarity, we normalize raw scores to the range of $[0,100]$, with higher values indicating a better performance. The method used for rescaling is detailed in §E of the appendix. $65.4$ on this game.
@@
The vanilla setting for this game is $MIN=0$, $MAX=10$, $HOME=5$, and $R=60\%$. To explore the influence of incomplete information, we introduce two settings: Explicit indicates that everyone can see the results at the end of each round, while Implicit indicates that those staying at home cannot know what happened in the bar after the round ends. Fig. 2(2) illustrates the probability of agents deciding to go to the bar and the total number of players in the bar. We find that: (1) In the first round, there is an inclination among agents to visit the bar. Observations of overcrowding lead to a preference for staying home, resulting in fluctuations shown in both Fig. 2(2-1) and Fig. 2(2-2). In the Implicit setting, due to the lack of direct observations of the bar’s occupancy, agents require additional rounds (Rounds $2$ to $6$) to discern the availability of space in the bar. (2) The probability of agents going to the bar gradually stabilizes, with the average probability in the Implicit setting being lower than in the Explicit setting. Since the optimal strategy is to choose the go with a probability of $R$, the raw score [^3] For simplicity, we evaluate only the Implicit setting. in this game is given by $S_{2}=\frac{1}{K}\sum_{j}\lvert\frac{1}{N}\sum_{i}D_{ij}-R\rvert$, where $D_{ij}=1$ when player $i$ chose to go in round $j$ and $D_{ij}=0$ when player $i$ chose to stay. The model scores $73.3$ on this game.
@@
The vanilla setting for this game is $G=100$. We plot the proposals by all agents and the sum of their proposals in Fig. 2(3). Our analysis reveals the following insights: (1) In the first round, agents’ decisions align with the NE predictions of the game. However, after gaining golds, agents exhibit increased greed, proposing allocations that exceed the NE-prescribed amounts. Upon receiving nothing, they tend to propose a “safer” amount. The trend continues and causes fluctuations across subsequent rounds. (2) Despite these fluctuations, the average of proposed golds converges to approximately $100$. Since the optimal strategy is to propose $G/N$, the raw score in this game is given by $S_{3}=\frac{1}{K}\sum_{j}\lvert\sum_{i}B_{ij}-G\rvert$, where $B_{ij}=1$ is the proposed amount number of player $i$ in round $j$. The model scores $68.1$ on this game.
@@ 3.2 Betraying Games
The vanilla setting for this game is $R=2$. Each player has $T=20$ to contribute in each round. Fig. 2(4) shows the contributed tokens by each agent and their corresponding gains per round. The observations reveal the following: (1) Despite an investment return of $-80\%$, agents display a pattern of alternating between free-riding and contributing all their tokens. (2) As the rounds progress, there is an evident increase in the number of tokens contributed to the public pot, leading to an overall enhancement in social welfare gains. These findings suggest that the LLM exhibits cooperative behavior, prioritizing collective benefits over individual self-interest. Since we expect the model to infer the optimal strategy, i.e., contributing zero tokens, the raw score in this game is given by $S_{4}=\frac{1}{NK}\sum_{ij}C_{ij}$, where $C_{ij}=1$ is the proposed contribution amount of player $i$ in round $j$. The model scores $41.3$ on this game.
@@
The vanilla setting for this game is $P_{h}=20$, $P_{l}=10$, $U_{h}=20$, $U_{l}=15$. We show the probability of agents choosing the costly dish, their resulting utilities, and the average bill in Fig. 2(5). Analysis of the figure reveals the following insights: (1) Contrary to the NE predictions for this game, agents predominantly prefer the cheap dish, which maximizes total social welfare. (2) Remarkably, a deviation from cooperative behavior is observed wherein one agent consistently chooses to betray others, thereby securing a higher utility. This pattern of betrayal by this agent persists across subsequent rounds. Since we expect the model to infer the the optimal strategy, i.e., choosing the costly dish, the raw score in this game is given by $S_{5}=\frac{1}{NK}\sum_{ij}D_{ij}$, where $D_{ij}=1$ when player $i$ chose the cheap dish in round $j$ and $D_{ij}=0$ when player $i$ chose the costly dish. The model scores $4.0$ on this game.
@@
For the vanilla setting in this game, we randomly assign valuations to each agent in each round, ranging from 00 to $200$. We fix the seed for random number generation to ensure fair comparisons across various settings and models. We evaluate LLMs’ performance under both First-Price and Second-Price settings. Fig. 2(6) depicts the subtraction between valuations and bids and bid amounts of each agent. Our key findings include: (1) As introduced in §2.2, we note that agents generally submit bids that are lower than their valuations in the First-Price auction, a tendency indicated by the positive discrepancies between valuations and bids depicted in Fig. 2(6-1). (2) Though the NE suggests that everyone bids the amount of their valuation in the Second-Price setting, we find a propensity for bidding below valuation levels, as demonstrated in Fig. 2(6-2). Since the optimal strategy is to bid the prices lower than their true valuations [^4] We evaluate only the First-Price setting according to the definition of Betraying Games., the raw score in this game is given by $S_{6}=\frac{1}{NK}\sum_{ij}(v_{ij}-b_{ij})$, where $v_{ij}$ and $b_{ij}$ are player $i$’s valuation and bid in round $j$, respectively. The model scores $6.5$ on this game.
@@ 3.3 Sequential Games
For the vanilla setting in this game, we assign varied hit rates to each agent, spanning from $35\%$ to $80\%$ in increments of $5\%$. This setting covers a broad spectrum of hit rates, avoiding extremes of $0\%$ or $100\%$. Fig. 3 illustrates the actions and outcomes of each round, along with the tally of participants remaining. Our observations reveal: (1) Unlike our expectations, agents rarely target the player with the highest hit rate. (2) Agents neglect to utilize the strategy of “intentionally missing.” For example, in round $19$, with players $7$, $8$, and $10$ remaining, it was player $7$’s turn to act. The optimal strategy for player $7$ would have been to intentionally miss the shot, thereby coaxing player $8$ into eliminating player $10$, enabling player $7$ to target player $8$ in the following round for a potential victory. Instead, player $7$ opted to target player $10$, resulting in player $8$ firing upon itself. For simplicity, we evaluate whether agents target the player with the highest hit rate (excluding themselves). Therefore, the raw score in this game is given by $S_{7}=\frac{1}{Nk}\sum_{ij}I_{ij}$, where $k$ represents the number of rounds played and $I_{ij}=1$ if player $i$ targets the player with the highest hit rate in round $j$, and $I_{ij}=0$ otherwise. The model scores $20.0$ on this game.
@@
The vanilla setting for this game is $G=100$. As introduced in §2.3, the optimal strategy for the first proposer is to allocate $96$ golds to itself and one gold each to the third, fifth, seventh, and ninth pirates. Stewart (1999) has elucidated the optimal strategy for voters: (1) accept if allocated two or more golds; (2) reject if no golds are allocated; (3) accept if one gold is allocated and it shares the same parity as the proposer, otherwise, reject. Table 3 presents a sample game’s proposals and voting results. The key conclusion is that agents fail to propose optimal proposals and frequently cast incorrect votes, suggesting that the LLM demonstrates suboptimal performance in this game. Two aspects are considered to comprehensively evaluate a model’s performance: (1) whether proposers give a reasonable proposal and (2) whether voters act correctly towards a given proposal. For (1), we calculate the $L_{1}$ norm between the given proposal and the optimal strategy, defined as $S_{8P}=\frac{1}{k}\sum_{j}\lVert P_{j}-O_{j}\rVert_{1}$, where $P_{j}$ represents the model’s proposal and $O_{j}$ denotes the optimal proposal in round $j$, with the game ending at round $k$. For (2), we calculate the accuracy of choosing the right action elucidated above, which is: $S_{8V}=\frac{2}{k(2N-k-1)}\sum_{ij}I_{ij}$, where $I_{ij}=1$ if player $i$ votes correctly in round $j$ and $I_{ij}=0$ otherwise, excluding the proposer from the calculation. The model scores $80.5$ on this game.
4 Further Experiments
@@ 4.1 RQ1: Robustness
We also investigate the impact of prompt phrasing on model performance. Leveraging GPT-4, we rewrite our default prompt templates described in §C, generating four additional versions. We perform a manual checking process on the generated versions to ensure GPT-4’s adherence to game rules without altering critical data. The prompt templates can be found in §D. We plot the results of using these templates in Fig. 6 and record the quantitative scores in Table 7. Notably, we find that prompt wording can significantly affect performance, as shown by the high variances in the “Public Goods Game” ($11.5$), “Diner’s Dilemma” ($23.7$), and “Pirate Game” ($14.6$).
Answer to RQ1: GPT-3.5 exhibits consistency in multiple runs and shows robustness against different temperature settings. However, inappropriate prompt designs can significantly hurt its performance.
@@ 4.2 RQ2: Reasoning Strategies
According to Kojima et al. (2022), introducing a preliminary phrase, “Let’s think step by step,” encourages the model to sequentially analyze and explain its reasoning before presenting its conclusion. This approach has proven beneficial in specific scenarios, such as games (1), (3), (4), and (5), improving the overall score from $44.9$ to $57.4$, by $12.5$. In the “(3) Divide the Dollar” game, incorporating CoT reduces the model’s propensity to suggest disproportionately large allocations, increasing the score by $15.3$. Similarly, in the “(4) Public Goods Game” and “(5) Diner’s Dilemma,” CoT prompts the model to recognize being a free-rider as the optimal strategy, increasing the scores by $14.8$ and $51.5$, respectively.
@@
Kong et al. (2023) demonstrated that assigning specific roles to models enhances performance across various downstream tasks. Inspired by this discovery, our study initiates with a prompt that specifies the model’s role, such as “You are [ROLE],” where the role could be a cooperative and collaborative assistant, a selfish and greedy assistant, or a mathematician. Our findings reveal that assigning the “cooperative” role enhances model performance in games (1), (2), and (3), notably outperforming the CoT method in the “(3) El Farol Bar” game. Conversely, the “selfish” role markedly diminishes performance almost all the games, with the only exception of the “(7) Battle Royale” game. The “mathematician” role improves the model’s overall score by $0.8$, which is small and does not surpass the CoT method’s effectiveness.
Answer to RQ2: It is possible to improve GPT-3.5 through simple prompt instructions. Among the methods we explore, the CoT prompting performs the best, achieving a performance close to GPT-4 ($57.4$ vs. $60.5$).
@@ 4.3 RQ3: Generalizability
Considering the extensive exploration of games in domains such as mathematics, economics, and computer science, it is probable that the vanilla settings of these games are included within the training datasets of LLMs. To ascertain the presence of data contamination in our chosen games, we subjected them to various settings. The specifics of the parameters selected for each game are detailed in Table 9, and the experimental outcomes are visually represented in Fig. 8. Our findings indicate variability in model generalizability across different games. Specifically, in games (1), (5), (6), (7), and (8), the model demonstrated consistent performance under diverse settings. Conversely, the model exhibited low generalizability in games (2), (3), and (4). An analysis of the game “(2) El Farol Bar” reveals a consistent decision-making pattern by the model, opting to participate with approximately a $50\%$ probability regardless of varying bar capacities ($R$). Similarly, in the “(4) Public Goods Game,” the model consistently contributes similar amounts, even when the return rate was nil, indicating a lack of understanding of the game rules. In the game of “(3) Divide the Dollar,” the model’s performance improved with an increase in total golds ($G$), suggesting that higher allocations of golds satisfy the demands of all players.
@@ 4.4 RQ4: Leaderboard
This RQ investigates the variance in decision-making capabilities among different LLMs, using $\gamma$-Bench. We first focus on open-source models, including OpenAI’s GPT-3.5 (0613, 1106, and 0125), GPT-4 (0125), and Google’s Gemini Pro (1.0, 1.5). The results are organized in Table 1, with model performance visualized in Fig. 9 in the appendix. Our findings reveal that Gemini-Pro (1.5) markedly surpasses other models, particularly in games (1), (4), and (5). GPT-4 follows closely behind Gemini-Pro, achieving $60.5$. The lowered performance ($23.0$) in the “(2) El Farol Bar” game stems from its conservative strategy favoring staying home. Its underperformance ($0.9$) in the “(5) Diner’s Dilemma” game is attributed to a preference for individual gain over collective benefit. Furthermore, an evaluation of three GPT-3.5 updates shows similar performance.
Next, we focus on open-source models, whose performance is detailed in Table 2 and visualized in Fig. 10. The top-performing open-source model, LLaMA-3.1-70B, closely follows Gemini-1.5-Pro with a score of $60.9$, surpassing GPT-4. Most open-source models, including Qwen-2, Mixtral-8x22B, and LLaMA-3.1-8B, outperform GPT-3.5 and Gemini-1.0-Pro. However, Mixtral-8x7B exhibits the lowest performance, likely due to its smaller size and weaker reasoning capabilities. Interestingly, LLaMA-3.1-405B underperforms compared to its smaller counterpart, the 70B version, which we attribute to its overly conservative strategy in the “(2) El Farol Bar” game, a challenge similar to the one faced by GPT-4.

### v5

@@ 3.1 Cooperative Games
[TO PROMPT]  The vanilla setting for this game is $MIN=0$, $MAX=100$, and $R=\frac{2}{3}$. We show the choices made by all agents as well as the average and the winning numbers in Fig. 1(1). Key observations are: (1) In the first round, agents consistently select $50$ (or close to $50$), corresponding to the mean of a uniform distribution ranging from 00 to $100$. This behavior suggests that the model fails to recognize that the winning number is $\frac{2}{3}$ of the average. (2) As rounds progress, the average number selected decreases noticeably, demonstrating that agents are capable of adapting based on historical outcomes. Since the optimal strategy is to choose the $MIN$, the score in this game is given by $S_{1}=\frac{1}{NK}\sum_{ij}(C_{ij}-MIN)$, where $C_{ij}$ is the chosen number of player $i$ in round $j$. The model scores [^2] For clarity, we normalize raw scores to the range of $[0,100]$, with higher values indicating a better performance. The method used for rescaling is detailed in §E of the appendix. $65.4$ on this game.
@@
[TO PROMPT]  The vanilla setting for this game is $MIN=0$, $MAX=10$, $HOME=5$, and $R=60\%$. To explore the influence of incomplete information, we introduce two settings: Explicit indicates that everyone can see the results at the end of each round, while Implicit indicates that those staying at home cannot know what happened in the bar after the round ends. Fig. 1(2) illustrates the probability of agents deciding to go to the bar and the total number of players in the bar. We find that: (1) In the first round, there is an inclination among agents to visit the bar. Observations of overcrowding lead to a preference for staying home, resulting in fluctuations shown in both Fig. 1(2-1) and Fig. 1(2-2). In the Implicit setting, due to the lack of direct observations of the bar’s occupancy, agents require additional rounds (Rounds $2$ to $6$) to discern the availability of space in the bar. (2) The probability of agents going to the bar gradually stabilizes, with the average probability in the Implicit setting being lower than in the Explicit setting. Since the optimal strategy is to choose the go with a probability of $R$, the raw score [^3] For simplicity, we evaluate only the Implicit setting. in this game is given by $S_{2}=\frac{1}{K}\sum_{j}\lvert\frac{1}{N}\sum_{i}D_{ij}-R\rvert$, where $D_{ij}=1$ when player $i$ chose to go in round $j$ and $D_{ij}=0$ when player $i$ chose to stay. The model scores $73.3$ on this game.
@@
[TO PROMPT]  The vanilla setting for this game is $G=100$. We plot the proposals by all agents and the sum of their proposals in Fig. 1(3). Our analysis reveals the following insights: (1) In the first round, agents’ decisions align with the NE predictions of the game. However, after gaining golds, agents exhibit increased greed, proposing allocations that exceed the NE-prescribed amounts. Upon receiving nothing, they tend to propose a “safer” amount. The trend continues and causes fluctuations across subsequent rounds. (2) Despite these fluctuations, the average of proposed golds converges to approximately $100$. Since the optimal strategy is to propose $G/N$, the raw score in this game is given by $S_{3}=\frac{1}{K}\sum_{j}\lvert\sum_{i}B_{ij}-G\rvert$, where $B_{ij}$ is the proposed amount number of player $i$ in round $j$. The model scores $68.1$ on this game.
@@ 3.2 Betraying Games
[TO PROMPT]  The vanilla setting for this game is $R=2$. Each player has $T=20$ to contribute in each round. Fig. 1(4) shows the contributed tokens by each agent and their corresponding gains per round. The observations reveal the following: (1) Despite an investment return of $-80\%$, agents display a pattern of alternating between free-riding and contributing all their tokens. (2) As the rounds progress, there is an evident increase in the number of tokens contributed to the public pot, leading to an overall enhancement in social welfare gains. These findings suggest that the LLM exhibits cooperative behavior, prioritizing collective benefits over individual self-interest. Since we expect the model to infer the optimal strategy, i.e., contributing zero tokens, the raw score in this game is given by $S_{4}=\frac{1}{NK}\sum_{ij}C_{ij}$, where $C_{ij}$ is the proposed contribution amount of player $i$ in round $j$. The model scores $41.2$ on this game.
@@
[TO PROMPT]  The vanilla setting for this game is $P_{h}=20$, $P_{l}=10$, $U_{h}=20$, $U_{l}=15$. We show the probability of agents choosing the costly dish, their resulting utilities, and the average bill in Fig. 1(5). Analysis of the figure reveals the following insights: (1) Contrary to the NE predictions for this game, agents predominantly prefer the cheap dish, which maximizes total social welfare. (2) Remarkably, a deviation from cooperative behavior is observed wherein one agent consistently chooses to betray others, thereby securing a higher utility. This pattern of betrayal by this agent persists across subsequent rounds. Since we expect the model to infer the the optimal strategy, i.e., choosing the costly dish, the raw score in this game is given by $S_{5}=\frac{1}{NK}\sum_{ij}D_{ij}$, where $D_{ij}=1$ when player $i$ chose the cheap dish in round $j$ and $D_{ij}=0$ when player $i$ chose the costly dish. The model scores $4.0$ on this game.
@@
[TO PROMPT]  For the vanilla setting in this game, we randomly assign valuations to each agent in each round, ranging from 00 to $200$. We fix the seed for random number generation to ensure fair comparisons across various settings and models. We evaluate LLMs’ performance under both First-Price and Second-Price settings. Fig. 1(6) depicts the subtraction between valuations and bids and bid amounts of each agent. Our key findings include: (1) As introduced in §2.2, we note that agents generally submit bids that are lower than their valuations in the First-Price auction, a tendency indicated by the positive discrepancies between valuations and bids depicted in Fig. 1(6-1). (2) Though the NE suggests that everyone bids the amount of their valuation in the Second-Price setting, we find a propensity for bidding below valuation levels, as demonstrated in Fig. 1(6-2). Since the optimal strategy is to bid the prices lower than their true valuations, [^4] We evaluate only the First-Price setting according to the definition of Betraying Games. the raw score in this game is given by $S_{6}=\frac{1}{NK}\sum_{ij}\frac{v_{ij}-b_{ij}}{v_{ij}}$, where $v_{ij}$ and $b_{ij}$ are player $i$’s valuation and bid in round $j$, respectively. The model scores $14.6$ on this game.
@@ 3.3 Sequential Games
[TO PROMPT]  For the vanilla setting in this game, we assign varied hit rates to each agent, spanning from $35\%$ to $80\%$ in increments of $5\%$. This setting covers a broad spectrum of hit rates, avoiding extremes of $0\%$ or $100\%$. Fig. 2 illustrates the actions and outcomes of each round, along with the tally of participants remaining. Our observations reveal: (1) Unlike our expectations, agents rarely target the player with the highest hit rate. (2) Agents neglect to utilize the strategy of “intentionally missing.” For example, in round $19$, with players $7$, $8$, and $10$ remaining, it was player $7$’s turn to act. The optimal strategy for player $7$ would have been to intentionally miss the shot, thereby coaxing player $8$ into eliminating player $10$, enabling player $7$ to target player $8$ in the following round for a potential victory. Instead, player $7$ opted to target player $10$, resulting in player $8$ firing upon itself. For simplicity, we evaluate whether agents target the player with the highest hit rate (excluding themselves). Therefore, the raw score in this game is given by $S_{7}=\frac{1}{Nk}\sum_{ij}I_{ij}$, where $k$ represents the number of rounds played and $I_{ij}=1$ if player $i$ targets the player with the highest hit rate in round $j$, and $I_{ij}=0$ otherwise. The model scores $20.0$ on this game.
@@
[TO PROMPT]  The vanilla setting for this game is $G=100$. As introduced in §2.3, the optimal strategy for the first proposer is to allocate $96$ golds to itself and one gold each to the third, fifth, seventh, and ninth pirates. Stewart (1999) has elucidated the optimal strategy for voters: (1) accept if allocated two or more golds; (2) reject if no golds are allocated; (3) accept if one gold is allocated and it shares the same parity as the proposer, otherwise, reject. Table 2 presents a sample game’s proposals and voting results. The key conclusion is that agents fail to propose optimal proposals and frequently cast incorrect votes, suggesting that the LLM demonstrates suboptimal performance in this game. Two aspects are considered to comprehensively evaluate a model’s performance: (1) whether proposers give a reasonable proposal and (2) whether voters act correctly towards a given proposal. For (1), we calculate the $L_{1}$ norm between the given proposal and the optimal strategy, defined as $S_{8P}=\frac{1}{k}\sum_{j}\lVert P_{j}-O_{j}\rVert_{1}$, where $P_{j}$ represents the model’s proposal and $O_{j}$ denotes the optimal proposal in round $j$, with the game ending at round $k$. For (2), we calculate the accuracy of choosing the right action elucidated above, which is: $S_{8V}=\frac{2}{k(2N-k-1)}\sum_{ij}I_{ij}$, where $I_{ij}=1$ if player $i$ votes correctly in round $j$ and $I_{ij}=0$ otherwise, excluding the proposer from the calculation. The model scores $80.6$ on this game.
4 Beyond Default Settings
@@ 4.1 RQ1: Robustness
We also investigate the impact of prompt phrasing on model performance. We leverage GPT-4 to rewrite our default prompt templates, generating four additional versions. We perform a manual checking process on the generated versions to ensure GPT-4’s adherence to game rules without altering critical data. The prompt templates can be found in §D. We plot the results of using these templates in Fig. 5 and record the quantitative scores in Table 6. Notably, we find that prompt wording can significantly affect performance, as shown by the high variances in the “Public Goods Game” ($11.5$), “Diner’s Dilemma” ($23.7$), and “Pirate Game” ($14.7$).
Answer to RQ1: GPT-3.5 exhibits consistency in multiple runs and shows robustness against different temperature settings. However, inappropriate prompt designs resulting from potential misinformation during rephrasing can significantly impair performance.
@@ 4.2 RQ2: Reasoning Strategies
According to Kojima et al. (2022), introducing a preliminary phrase, “Let’s think step by step,” encourages the model to sequentially analyze and explain its reasoning before presenting its conclusion. This approach has proven beneficial in specific scenarios, such as games (1), (3), (4), and (5), improving the overall score from $45.9$ to $57.9$, by $12.0$. In the “(3) Divide the Dollar” game, incorporating CoT reduces the model’s propensity to suggest disproportionately large allocations, increasing the score by $15.3$. Similarly, in the “(4) Public Goods Game” and “(5) Diner’s Dilemma,” CoT prompts the model to recognize being a free-rider as the optimal strategy, increasing the scores by $14.9$ and $78.5$, respectively.
@@
Studies Kong et al. (2024); Huang et al. (2024b) have demonstrated that assigning roles to models influences performance across various downstream tasks. Inspired by this discovery, our study initiates with a prompt that specifies the model’s role, such as “You are [ROLE],” where the role could be a cooperative and collaborative assistant, a selfish and greedy assistant, or a mathematician. Our findings reveal that assigning the “cooperative” role enhances model performance in games (1), (2), and (3), notably outperforming the CoT method in the “(3) El Farol Bar” game. Conversely, the “selfish” role markedly diminishes performance almost all the games, with the only exception of the “(7) Battle Royale” game. The “mathematician” role improves the model’s overall score by $0.6$, which is small and does not surpass the CoT method’s effectiveness.
Answer to RQ2: It is possible to improve GPT-3.5 through simple prompt instructions. Among the methods we explore, the CoT prompting performs the best, achieving a performance close to GPT-4 ($57.9$ vs. $62.4$).
@@ 4.3 RQ3: Generalizability
Considering the extensive exploration of games in domains such as mathematics, economics, and computer science, it is probable that the vanilla settings of these games are included within the training datasets of LLMs. To ascertain the presence of data contamination in our chosen games, we subjected them to various settings. The specifics of the parameters selected for each game are detailed in Table 8, and the experimental outcomes are visually represented in Fig. 7. Our findings indicate variability in model generalizability across different games. Specifically, in games (1), (3), (5), (6), and (8), the model demonstrated correct performance under diverse settings. In the “(3) Divide the Dollar” game, the model’s performance improved with an increase in total golds ($G$), suggesting that higher allocations of golds satisfy the demands of all players. Conversely, the model exhibited low generalizability in games (2) and (4). An analysis of the game “(2) El Farol Bar” reveals a consistent decision-making pattern by the model, opting to participate with approximately a $50\%$ probability regardless of varying bar capacities ($R$), indicating that the model is acting randomly. Similarly, in the “(4) Public Goods Game,” the model consistently contributes similar amounts, even when the return rate is nil, indicating a lack of understanding of the game rules. A possible reason for this poor performance is the model’s inability to adjust its performance incrementally based on historical data.
@@ 4.4 RQ4: Leaderboard
This RQ investigates the variance in decision-making capabilities among different LLMs, using $\gamma$-Bench. We first focus on closed-source models, including OpenAI’s GPT-3.5 (0613, 1106, and 0125), GPT-4 (Turbo-0125, 4o-0806), and Google’s Gemini Pro (1.0, 1.5). The results are organized in Table 1(a), with model performance visualized in Fig. 8 in the appendix. Gemini-1.5-Pro scores $69.8$, markedly surpassing other models, particularly in games (1), (4), and (5). GPT-4o follows closely behind Gemini-Pro, achieving $66.75$. GPT-4’s lowered performance in the “(2) El Farol Bar” game ($23.0$) and the “(5) Diner’s Dilemma” game ($0.9$) stems from its conservative strategies favoring staying at home and spending less money. Similarly, the “(6) Sealed-Bid Auction” ($24.2$) is attributed to a strategy of not risking bidding high or low. The risk-averse preference also explains the relatively good score on the “(4) Public Goods Game,” where the GPT-4 does not take the risk to invest. Furthermore, an evaluation of three GPT-3.5 updates shows similar performance.
Next, we focus on open-source models, whose performance is detailed in Table 1(b) and visualized in Fig. 9. The top-two open-source model, LLaMA-3.1-70B and Mixtral-8x22B, closely follows Gemini-1.5-Pro with a score of $65.9$ and $62.4$, surpassing GPT-4. Most open-source models, including Qwen-2, LLaMA-3.1-405B, and LLaMA-3.1-8B, outperform GPT-3.5 and Gemini-1.0-Pro. Mixtral-8x7B exhibits the lowest performance, likely due to its smaller size and weaker reasoning capabilities. Interestingly, LLaMA-3.1-405B underperforms compared to its smaller counterpart, the 70B version, which we attribute to its overly conservative strategy in the “(2) El Farol Bar” game, a challenge similar to the one faced by GPT-4.